# Reinforcement learning for anisotropic p-adaptation and error estimation in high-order solvers

### v1

@@ 3.4.4 Training
The reinforcement learning agent is trained to select the optimal polynomial order for each element of the computational mesh. The proposed strategy for the training process does not depend on a CFD solver. This decision results in a faster and general training that does not depend on the problem to be solved. As explained in Section 3.3, we select a value iteration algorithm, which is a classic model-based reinforcement learning method. The training process consists in executing an iterative algorithm until convergence is reached (details of the algorithm are shown in C). The v-values are initialized to $v_{0}=0$ and the value function for each state is calculated using the update rule from eq. (3). In this equation, the discount factor $\gamma=0.5$ is known and fixed. This low value for $\gamma$ has been selected to increase the importance of the short-term, as it was explained in the previous Section 3.4.3. The remaining variables in the right hand side of the update rule are the probabilities $\mathbb{P}(s^{\prime},r\mid s,a)$, the reward function $r$ and the future states $s^{\prime}$. All of them require additional considerations to be computed. In particular, a reference solution, which plays the role of the analytical solution of the problem during the training, has to be defined, together with the polynomial approximation for each state within the DGSEM framework.
Definition of the approximated solution from the state vector
For each state $s$, we only know the values of the approximated solution at the Gauss nodes and the polynomial order $p=size(s)-1$. This information allows to reconstruct, for each state, the approximated solution $y^{*}$ within the DGSEM framework, which is generated through Lagrange polynomials based on the Gauss nodes:
@@
$\ell_{j}(x)=\prod_{\begin{subarray}{c}i=0\\ i\neq j\end{subarray}}^{k}\frac{x-x_{i}}{x_{j}-x_{i}}=\frac{x-x_{0}}{x_{j}-x_{0}}\cdots\frac{x-x_{j-1}}{x_{j}-x_{j-1}}\frac{x-x_{j+1}}{x_{j}-x_{j+1}}\cdots\frac{x-x_{k}}{x_{j}-x_{k}},$ (9)
@@
The reference solution, $y$, represents possible outcomes that can be found within one element in a real simulation. In particular, we define the reference solution as a polynomial function. This hypothesis is reasonable within the discontinuous Galerkin framework, as the final solution is approximated using Lagrange polynomials (see Section 3.1). Therefore, we limit the space of possible reference solutions to polynomials of different orders. For each state, $s$, we consider three possible scenarios, where the reference solution, $y$, is a polynomial function:
@@
    One order higher than the approximated solution $y^{*}$.
@@
    Of the same order as the approximated solution $y^{*}$.
@@
    One order lower than the approximated solution $y^{*}$.
@@
    In the second case, the reference solution, $y$, is exactly the approximated solution, $y^{*}$.
@@
    Finally, in the third case it is necessary to check whether it is possible to approximate the current state with a low-order polynomial. Hence, the approximated solution for the present state, $y^{*}$, is evaluated at the low-order Gauss nodes; and then, those values are used to define a low-order polynomial using eq. (8) and (9). Then, the low-order (possible reference solution) and the current polynomial (approximated solution) are compared using the same root mean squared error defined for the reward in eq. (6). This step is sketched in Figure 4. If the $rmse$ is below a threshold $rmse_{th}=0.1$, then we consider that the low-order reference solution is possible for the current state. In other case, we reject this third solution.
Figure 4: Example of a comparison between the approximated solution $y^{*}$ ($p=3$ in this example) and the low-order reference solution ($p=2$ in this example). The error between both functions is estimated through the root mean squared error defined in eq. (6).
@@
The probabilities $\mathbb{P}(s^{\prime},r\mid s,a)$ define the likelihood of being in one or another scenario; that is, the likelihood of using one reference solution over the other ones. In this case, since the real analytical solution will strongly depend on each specific problem, we consider that all reference solutions are equally likely. Therefore, when all three scenarios are possible, $\mathbb{P}(s^{\prime},r\mid s,a)=1/3$ However, when the third scenario (i.e. the reference solution is a polynomial function one order lower than the approximated solution) is not possible (the current state cannot be well captured using a low-order polynomial function), then $\mathbb{P}(s^{\prime},r\mid s,a)=1/2$ for the first two scenarios, and $\mathbb{P}(s^{\prime},r\mid s,a)=0$ for the third scenario.
@@
(a) Example of values at the Gauss nodes for a $p=3$ approximated solution.
@@
Figure 5: a) Values of the state vector for an approximated solution $y^{*}(x)$ and b) reference solutions $y(x)$ computed during the training process.
For each scenario, $j=1,\,2\,\textrm{and}\,3$, we calculate the reward $r_{j}(s)$ using eq. (7). Then, for each action $a$ , we compute the next state $s^{\prime}_{j,a}$ by evaluating the reference solution of the current scenario, $j$, into the Gauss nodes for the new polynomial order $p^{\prime}=p+a$. It is important to note that the new polynomial order, $p^{\prime}$, must be bounded between $p_{\min}$ and $p_{\max}$. Finally, once we have computed the reference solutions, the rewards and the next states for each scenario, we have to introduce all the values in the value iteration algorithm. Using the new notation for each scenario:
@@
Both rules drastically reduce the number of possible states during the training phase, and hence the effective training time. Overall, the computational complexity of the value iteration algorithm grows exponentially with the maximum polynomial order considered. Furthermore, since the number of states shows the same trend as the computational complexity, there is a corresponding increase in the memory requirements needed to store all the v-values. Despite these challenges, the algorithm is highly parallelizable, which enhances its scalability even with high polynomial orders. More importantly, training is only done once and offline. The same trained agent then is used for multiple cases (all the results provided in Section 4 are simulated with the same trained agent).
@@ 3.4.5 Error Estimation
One may notice that, although we estimate the reward $\hat{r}$ two steps in the future, the polynomial order used in eq. (17) is the polynomial order $p$ of the present state $s$. The reason behind this decision is that we want to estimate the error that we potentially have based on the measurements from the environment that can be collected in real time. Therefore, the use of the same polynomial order as is used in the environment when computing the state leads to more accurate estimations.
It is important to note that this error estimation is the error that the RL agent believes to be real, but the only feedback received by the agent comes from the reward function. In addition, the average rewards $\bar{r}$ and $\bar{r}^{\prime}$ are calculated based on the probabilities defined during the training. Hence, it may be impossible to obtain an estimate for some specific situations. However, even if an estimation of the error cannot be obtained, the solution has a true meaning behind it.
@@
    $v^{*}(s)-\bar{r}-\gamma\bar{r}^{\prime}<0$: The agent believes that the error is much higher than expected. This situation is unusual, but it may happen if the polynomial order used to solve the problem is very small compared to the optimum polynomial order (e.g. $p=2$ and $p^{*}>4$).
@@
    $v^{*}(s)-\bar{r}-\gamma\bar{r}^{\prime}>1$: The agent believes that the error is very small compared to the standard deviation of the reward $\sigma$ (see eq. (7) for details). This situation may happen if the polynomial order is high in relation to the optimum polynomial order (e.g. $p=4$ and $p^{*}\leq 2$).
@@
This error estimation will be tested in the results section.
@@ 3.4.6 Error estimation extrapolation for 3D problems
As we can use the RL p-adaptation agent for 3D problems, though it was trained with 1D cases (see Section 3.4.2), we also want to apply a similar approach to use the error estimation capability for 3D problems. In a similar way as we did for the state, inside each element we compute the estimated error for each row of Gauss nodes that aligns with each local axis $x$, $y$, $z$, and we average the results among every row of nodes, resulting in three different values of the error, one for each direction. Then, depending on how these three values are combined, we may obtain different error estimations for 3D problems. In particular, we provide two possible alternatives:
@@
    Average error: $\widehat{rmse_{q}}=\frac{\widehat{rmse_{q,x}}+\widehat{rmse_{q,y}}+\widehat{rmse_{q,z}}}{3}$. This approach provides, in general, a good estimation of the error, which is of the same order of magnitude than the real error for the selected variable $q$.
@@
    Maximum error: $\widehat{rmse_{q}}=\max{\left(\widehat{rmse_{q,x}},\,\widehat{rmse_{q,y}},\,\widehat{rmse_{q,z}}\right)}$. This estimation of the error tends to slightly overestimate the real error. Furthermore, it could be use as a maximum value that should not be reached in the real solution.
@@ 3.5 Implementation of p-adaptation in DGSEM solvers
First, we have to consider that the p-adaptation process in a CFD solver is highly time-consuming. Although the time required to evaluate the RL agent is negligible, as it is based on a look-up table, the whole adaptation process requires significant memory reallocation and, in some cases, dynamic load balancing in MPI simulations. Consequently, adapting the mesh at every iteration is impractical and computationally prohibitive. Therefore, we should adapt the mesh every $100-10000$ iterations, or when we have detected that it is really necessary to perform an adaptation (e.g. based on the error estimation). Furthermore, a sharp increase or decrease in the polynomial order may lead to a numerical divergence. For that reason, each time the mesh is adapted, only one action is taken in each element; that is, the polynomial order can be increased or decreased by 1 unit.
@@
Furthermore, when a new simulation is started, the original mesh is usually not p-adapted. An usual practice is to initialize all elements with a uniform low polynomial (e.g. $p=2$) to reduce the computational cost during the first stage of the simulation, which is, in general, a transient phase of little interest. However, this is not a fixed rule: the simulation can also be started from an uniform high-order polynomial, or even from another p-adapted mesh. Later, the p-adaptation algorithm may take control over the simulation and decide which polynomial order is desirable in each region.
@@
Additionally, a common problem that may arise in a real implementation is related to the state (set of scaled values at the Gauss nodes) that is provided to the RL agent to perform the adaptation, which is always scaled between $[-1,1]$, as previously explained in Section 3.4.1. Although this step is completely necessary to obtain a RL agent as general as possible, it may be problematic if the values at the Gauss nodes that allow us to create the state are extremely small. In those cases, the variable of interest that has been measured at the nodes can be considered to be constant along that row of points. However, after the scaling process, the state is not constant anymore, and the RL agent may perceive solution gradients that are actually negligible. To solve this issue, when the spatial change inside one element in the non-dimensional variable used to compute the state, $q$, is below a small tolerance, $\Delta q<5\cdot 10^{-3}$, we set the state to 0. By doing so, we can provide this additional information to the RL agent, obtaining a significant improvement in the performance. In addition, as the variable of interest $q$ is, in our solver, always nondimensional, this tolerance is general enough to be used in a wide variety of problems without modifying its default value. The default tolerance has been successfully used for all the problems reported in this work and for the different reported variables. Finally, note that tolerance is independent of the training process, and hence the RL agent, previously trained, is still valid for any value of the aforementioned tolerance.
@@
Another implementation decision is that the polynomial adaptation is bounded between $p_{\min}=2$ and $p_{\max}=6$. This range is appropriate for a wide variety of problems; however, in large simulations it could be desirable to decrease the polynomial below $p_{\min}=2$ in some regions, reducing the computational cost. To do so, we consider that if nothing relevant is happening inside one element; that is, the maximum change in the variable of interest is below the aforementioned tolerance, $\Delta q$, then we can decrease the polynomial order to $p=1$. Once we detect that the variations of the variable are above the tolerance, we switch to $p=2$ and the RL agent is in charge of the adaptation again.
@@
Also, it is important to decide which variables of interest (it can be more than one) should be considered to perform the adaptation process. For all the following problems, the state will be computed based on three non-dimensional components of momentum:
@@ 4 Results
To show the applicability and flexibility of the proposed methodology, different problems with increasing complexity are included. For all the cases, we show the effectiveness of the RL-based p-adaptation, leading to accurate solutions and a significant reduction of the computational time, when compared to using homogeneous polynomials and other adaptation criteria. Note that for those problems in which the error estimation methodology is used, we consider the average error as default (see Section 3.4.6), unless explicitly stated otherwise.
@@ 4.1 Error estimation for an Euler flow around a cylinder
The error estimation and the real error between the reference and the RL p-adapted solution, for the non-dimensional $u$ velocity, are shown in Figure 9. The error estimator is capable of providing the trend of the error, highlighting those regions where the real error is higher. Of course the exact value of the error cannot be estimated, given that the RL agent is exactly the same for every problem and PDE and no information regarding the equations to be solved, the mesh or the boundary conditions are provided. Even with this little knowledge, the RL agent shows an excellent behavior and versatility, and a fairly good estimation can be obtained.
@@
Figure 9: Comparison of the error estimation from the RL agent and the real error for the non-dimensional $u$ velocity under steady-state conditions for the Euler flow around a cylinder at $M=0.1$. The error was computed by comparing the p-adapted solution with a reference solution uniformly set to $p=5$. Only those elements of the mesh with an error higher than $10^{-6}$ are represented.

### v2

@@ 3.4.4 Training
The reinforcement learning agent is trained to select the optimal polynomial order for each element of the computational mesh. The proposed strategy for the training process does not depend on a CFD solver. This decision results in a faster and general training that does not depend on the problem to be solved. As explained in Section 3.3, we select a value iteration algorithm, which is a classic model-based reinforcement learning method. The training process consists in executing an iterative algorithm until convergence is reached (details of the algorithm are shown in C). The v-values are initialized to $v_{0}=0$ and the value function for each state is calculated using the update rule from eq. (3). In this equation, the discount factor $\gamma=0.5$ is known and fixed. This low value for $\gamma$ has been selected to increase the importance of the short term, as explained in the previous Section 3.4.3. The remaining variables on the right-hand side of the update rule are the probabilities $\mathbb{P}(s^{\prime},r\mid s,a)$, the reward function $r$ and the future states $s^{\prime}$. All of them require additional considerations to be computed. In particular, a reference solution, which plays the role of the analytical solution of the problem during the training, has to be defined, together with the polynomial approximation for each state within the DGSEM framework.
Definition of the approximate solution from the state vector
For each state $s$, we only know the values of the approximate solution at the Gauss nodes and the polynomial order $p=size(s)-1$. This information allows to reconstruct, for each state, the approximate solution $y^{*}$ within the DGSEM framework, which is generated through Lagrange polynomials based on the Gauss nodes:
@@
$\ell_{j}(x)=\prod_{\begin{subarray}{c}i=0\\ i\neq j\end{subarray}}^{k}\frac{x-x_{i}}{x_{j}-x_{i}}$ (9)
@@
The reference solution, $y$, represents possible outcomes that can be found within an element in a real simulation. In particular, we define the reference solution as a polynomial function. This hypothesis is reasonable within the discontinuous Galerkin framework, as the final solution is approximated using Lagrange polynomials (see Section 3.1). Therefore, we limit the space of possible reference solutions to polynomials of different orders. For each state, $s$, we consider three possible scenarios, where the reference solution, $y$, is a polynomial function:
@@
    One order higher than the approximate solution $y^{*}$.
@@
    Of the same order as the approximate solution $y^{*}$.
@@
    One order lower than the approximate solution $y^{*}$.
@@
    In the second case, the reference solution, $y$, is exactly the approximate solution, $y^{*}$.
@@
    Finally, in the third case it is necessary to check whether it is possible to approximate the current state with a low-order polynomial. Hence, the approximate solution for the present state, $y^{*}$, is evaluated at the low-order Gauss nodes; and then, these values are used to define a low-order polynomial using eq. (8) and (9). Next, the low-order (possible reference solution) and the current polynomial (approximate solution) are compared using the same root mean squared error defined for the reward in eq. (6). This step is sketched in Figure 4. If the $rmse$ is below a threshold $rmse_{th}=0.1$, then we consider that the low-order reference solution is possible for the current state. In other case, we reject this third solution.
Figure 4: Example of a comparison between the approximate solution $y^{*}$ ($p=3$ in this example) and the low-order reference solution ($p=2$ in this example). The error between both functions is estimated through the root mean squared error defined in eq. (6).
@@
The probabilities $\mathbb{P}(s^{\prime},r\mid s,a)$ define the likelihood of being in one or another scenario; that is, the likelihood of using one reference solution over the other ones. In this case, since the real analytical solution will strongly depend on each specific problem, we consider that all reference solutions are equally likely. Therefore, when all three scenarios are possible, $\mathbb{P}(s^{\prime},r\mid s,a)=1/3$ However, when the third scenario (i.e., the reference solution is a polynomial function one order lower than the approximate solution) is not possible (the current state cannot be well captured using a low-order polynomial function), then $\mathbb{P}(s^{\prime},r\mid s,a)=1/2$ for the first two scenarios, and $\mathbb{P}(s^{\prime},r\mid s,a)=0$ for the third scenario.
@@
(a) Example of values at the Gauss nodes for a $p=3$ approximate solution.
@@
Figure 5: a) Values of the state vector for an approximate solution $y^{*}(x)$ and b) reference solutions $y(x)$ calculated during the training process.
For each scenario, $j=1,\,2\,\textrm{and}\,3$, we calculate the reward $r_{j}(s)$ using eq. (7). Then, for each action $a$ , we compute the next state $s^{\prime}_{j,a}$ by evaluating the reference solution of the current scenario, $j$, into the Gauss nodes for the new polynomial order $p^{\prime}=p+a$. It is important to note that the new polynomial order, $p^{\prime}$, must be bounded between $p_{\min}$ and $p_{\max}$. Finally, once we have computed the reference solutions, the rewards, and the next states for each scenario, we have to introduce all the values in the value iteration algorithm. Using the new notation for each scenario:
@@
Both rules drastically reduce the number of possible states during the training phase, and hence the effective training time. Overall, the computational complexity of the value iteration algorithm grows exponentially with the maximum polynomial order considered. Furthermore, since the number of states shows the same trend as the computational complexity, there is a corresponding increase in the memory requirements needed to store all the v-values. Despite these challenges, the algorithm is highly parallelizable, which enhances its scalability even with high polynomial orders. More importantly, training is only done once and offline. The same trained agent is then used for multiple cases (all the results provided in Section 4 are simulated with the same trained agent).
@@ 3.4.5 Error Estimation
One may notice that, although we estimate the reward $\hat{r}$ in two steps in the future, the polynomial order used in eq. (17) is the polynomial order $p$ of the present state $s$. The reason behind this decision is that we want to estimate the error that we potentially have based on the measurements from the environment that can be collected in real time. Therefore, the use of the same polynomial order as is used in the environment when computing the state leads to more accurate estimations.
It is important to note that this error estimation is the error that the RL agent believes to be real, but the only feedback received by the agent comes from the reward function. In addition, the average rewards $\bar{r}$ and $\bar{r}^{\prime}$ are calculated based on the probabilities defined during training. Hence, it may be impossible to obtain an estimate for some specific situations. However, even if an estimate of the error cannot be obtained, the solution has a true meaning behind it.
@@
    $v^{*}(s)-\bar{r}-\gamma\bar{r}^{\prime}<0$: The agent believes that the error is much higher than expected. This situation is unusual, but it may happen if the polynomial order used to solve the problem is very small compared to the optimal polynomial order (e.g., $p=2$ and $p^{*}>4$).
@@
    $v^{*}(s)-\bar{r}-\gamma\bar{r}^{\prime}>1$: The agent believes that the error is very small compared to the standard deviation of the reward $\sigma$ (see eq. (7) for details). This situation may occur if the polynomial order is high in relation to the optimum polynomial order (e.g., $p=4$ and $p^{*}\leq 2$).
@@
This error estimation can be computed beforehand during the training process, with the exception of the scaling factor for the specific variable $q$. Therefore, the computational cost of this estimator at run time, in a real simulation, is negligible, resulting in an inexpensive error estimator that can be applied to any problem or variable of interest. The effectiveness of the error estimator will be tested in the results section.
@@ 3.4.6 Error estimation extrapolation for 3D problems
As we can use the RL p-adaptation agent for 3D problems, though it was trained with 1D cases (see Section 3.4.2), we also want to apply a similar approach to use the error estimation capability for 3D problems. In a similar way as we did for the state, inside each element we compute the estimated error for each row of Gauss nodes that aligns with each local axis $x$, $y$ and $z$, and we average the results among every row of nodes, resulting in three different values of the error, one for each direction. Then, depending on how these three values are combined, we may obtain different error estimations for 3D problems. In particular, we provide two possible alternatives:
@@
    Average error: $\widehat{rmse_{q}}=\frac{\widehat{rmse_{q,x}}+\widehat{rmse_{q,y}}+\widehat{rmse_{q,z}}}{3}$. This approach provides, in general, a good estimate of the error, which is of the same order of magnitude as the real error for the selected variable $q$.
@@
    Maximum error: $\widehat{rmse_{q}}=\max{\left(\widehat{rmse_{q,x}},\,\widehat{rmse_{q,y}},\,\widehat{rmse_{q,z}}\right)}$. This estimate of the error tends to slightly overestimate the real error. Furthermore, it could be used as a maximum value that should not be reached in the real solution.
The choice of one alternative or another to accurately estimate the error depends on each specific problem. Both methods will be compared in the results section.
@@ 3.5 Implementation of p-adaptation in DGSEM solvers
First, we have to consider that the p-adaptation process in a CFD solver is highly time-consuming. Although the time required to evaluate the RL agent is negligible, as it is based on a look-up table, the whole adaptation process requires significant memory reallocation and, in some cases, dynamic load balancing in MPI simulations. Consequently, adapting the mesh at every iteration is impractical and computationally prohibitive. Therefore, we should adapt the mesh every $100-10000$ iterations, or when we have detected that it is really necessary to perform an adaptation (e.g., based on the RL-based error estimation). Furthermore, a sharp increase or decrease in the polynomial order may lead to a numerical divergence. For that reason, each time the mesh is adapted, only one action is taken in each element; that is, the polynomial order can be increased or decreased by 1 unit.
@@
Furthermore, when a new simulation is started, the original mesh is usually not p-adapted. A common practice is to initialize all elements with a uniform low polynomial (e.g., $p=2$) to reduce computational costs during the first stage of the simulation, which is, in general, a transient phase of little interest. However, this is not a fixed rule: the simulation can also be started from an uniform high-order polynomial, or even from another p-adapted mesh. Later, the p-adaptation algorithm may take control over the simulation and decide which polynomial order is desirable in each region.
@@
Additionally, a common problem that may arise in a real implementation is related to the state (set of scaled values at the Gauss nodes) that is provided to the RL agent to perform the adaptation, which is always scaled between $[-1,1]$, as previously explained in Section 3.4.1. Although this step is completely necessary to obtain an RL agent as general as possible, it may be problematic if the values at the Gauss nodes that allow us to create the state are extremely small. In those cases, the variable of interest that has been measured at the nodes can be considered to be constant along that row of points. However, after the scaling process, the state is no longer constant and the RL agent may perceive solution gradients that are actually negligible. To solve this issue, when the spatial change inside one element in the non-dimensional variable used to compute the state, $q$, is below a small tolerance, $\Delta q<5\cdot 10^{-3}$, we set the state to 0. By doing so, we can provide this additional information to the RL agent, obtaining a significant improvement in performance. In addition, as the variable of interest $q$ is, in our solver, always non-dimensional, this tolerance is general enough to be used in a wide variety of problems without modifying its default value. The default tolerance has been successfully used for all the problems reported in this work and for the different reported variables. Finally, note that the tolerance is independent of the training process, and hence the RL agent, previously trained, is still valid for any value of the aforementioned tolerance.
@@
Another implementation decision is that the polynomial adaptation is bounded between $p_{\min}=2$ and $p_{\max}=6$. This range is appropriate for a wide variety of problems; however, in large simulations, it could be desirable to decrease the polynomial below $p_{\min}=2$ in some regions, reducing the computational cost. To do so, we consider that if nothing relevant is happening inside one element; that is, the maximum change in the variable of interest is below the aforementioned tolerance, $\Delta q$, then we can decrease the polynomial order to $p=1$. Once we detect that the variations of the variable are above the tolerance, we switch to $p=2$ and the RL agent is in charge of the adaptation again.
@@
In addition, it is important to decide which variables of interest (there can be more than one) should be considered to perform the adaptation process. For all the following problems, the state will be computed based on three non-dimensional components of momentum:
@@ 4 Results
To show the applicability and flexibility of the proposed methodology, different problems with increasing complexity are included. For all the cases, we show the effectiveness of the RL-based p-adaptation, leading to accurate solutions and a significant reduction of the computational time, when compared to using homogeneous polynomials and other adaptation criteria. Note that for those problems in which the RL-based error estimation methodology is used, we consider the average error as default (see Section 3.4.6), unless explicitly stated otherwise.
@@ 4.1 Error estimation for an Euler flow around a cylinder
The average error estimation and the real error between the reference and the RL p-adapted solution, for the non-dimensional $u$ velocity, are shown in Figure 9. The error estimator is capable of providing the trend of the error, highlighting those regions where the real error is higher. Of course, the exact value of the error cannot be estimated, given that the RL agent is exactly the same for every problem and PDE and no information regarding the equations to be solved, the mesh or the boundary conditions is provided. Even with this little knowledge, the RL agent shows an excellent behavior and versatility, and a fairly good estimation can be obtained. Furthermore, the computational cost of the estimation is negligible, as the error estimator is (almost) completely defined during the training of the RL agent.
@@
Figure 9: Comparison of the error estimated by the RL agent and the real error for the non-dimensional $u$ velocity under steady-state conditions for the Euler flow around a cylinder at $M=0.1$. The error was computed by comparing the p-adapted solution with a uniform $p=5$ reference solution. Only those elements of the mesh with an error higher than $10^{-6}$ are represented.